OpenStudy (anonymous):

Help on algebra 2 questions!!!! :( what are the roots of y=(x-3)^2-5

OpenStudy (anonymous):

Roots also means zeroes. SO you're trying to find the zeros of this equation. What does x equal.

OpenStudy (anonymous):

Hold on I lost it. :/

OpenStudy (anonymous):

is it 5?

OpenStudy (anonymous):

First make y=0

OpenStudy (anonymous):

could you make the equation into 0=(x-3)(x-3)-5 ??

OpenStudy (anonymous):

You could, but it's easier in the other form.\[0=(x-3)^{2}+5\] then subtract 5 on both sides

OpenStudy (anonymous):

sorry its supposed to be subtract 5, so add 5 on both sides \[5=(x-3)^{2}\]

OpenStudy (anonymous):

ok so then would you square root it ?

OpenStudy (anonymous):

YES

OpenStudy (anonymous):

2.23= (x-3)?

OpenStudy (anonymous):

Okay sure, but remember when you have a square root like the one on 5. Remember it has an invisible +/- sign

OpenStudy (anonymous):

\[x-3=\pm \sqrt{5}\]

OpenStudy (anonymous):

so then you add 3 ? so the answer can be negative and positive?

OpenStudy (anonymous):

Because remember the fact that the equation has a degree of 2. SO you have 2 zeroes.

OpenStudy (anonymous):

Well when do \[3-\sqrt{5}\] it's still positive.

OpenStudy (anonymous):

Some teachers want you to stop there. or subtract

OpenStudy (anonymous):

If they want you to the you will get x=0.76 and x=5.23

OpenStudy (anonymous):

But because you estimated the decimals. It isn't exact. SO i would leave it in the square root form.

OpenStudy (anonymous):

OOHHHH! ok!!!!!! thank you!!!!!! :)

OpenStudy (anonymous):

You're very welcome.